# Mapping Still Matters: Coarse-Graining with Machine Learning Potentials

### v1

@@ 1 Introduction
In this study, we investigate how modern E(3)-equivariant machine learning potentials can be applied to coarse-graining without the use of any priors. Using the MACE architecture as a representative modelMACE, we perform a systematic study to evaluate how the choice of mapping, species encoding and model parameters influence the stability and learned representation of the model. We compare these results with a classical CG potential to assess whether the learned models provide a more accurate description of coarse-grained energy landscapes. We evaluated three distinct systems of increasing complexity: liquid hexane, single amino acids, and a 15-mer polyalanine. We validate key findings by testing them with NequIP batzner20223.
@@ 2.1 Coarse-graining
In coarse-graining (CG), the degrees of freedom of an atomistic system are reduced by grouping different atoms into beads. Here we limit the study to mappings, in which each atom at most contributes to one CG bead, and use the center-of-mass (COM) mapping to derive CG positions and forces. Other, more general mapping choicesbabadi2006coarse, gay1981modification, Voth2008-zb, as well as optimal mappingsdiggins2018optimal, yang2023slicing, wang2019coarse, have also been explored, but remain less prevalent in practice. In the COM mapping, the CG positions $\textbf{R}\in\mathcal{R}^{3N}$ are derived as a linear mapping M of the atomistic positions $\textbf{r}\in\mathcal{R}^{3n}$, where $N<n$:
@@ 2.3 Coarse-grained Potential
To parameterize an equivariant MLP $U_{MLP}(\mathbf{R};\theta)$, we employ the MACE architectureMACE. MACE combines the idea of atomic cluster expansiondrautz2019atomic with message-passing neural networks (MPNNs)gilmer2017neural. In MACE, the messages passed between nodes, in our case CG beads, are expanded using the idea of a hierarchical body order expansion
@@ 2.7 Well-tempered Metadynamics
where $h$ is the initial Gaussian height, $\sigma_{i}$ is the width of the $i$-th collective variable, and $\gamma=T^{*}/T$ is the bias factor. This corresponds to sampling the collective variables at an effective temperature $T^{*}=\gamma T$, which scales the free energy barriers by $1/\gamma$.
@@ 3.1 Liquid Hexane
Table 1: Comparison of a 2-layer MACE model with different correlation orders $\nu$ and average number of neighbors $N_{avg}$. Results are based on 50 × 1000 ps simulations. Green highlighting indicates better performance (arrows indicate direction: $\uparrow$ higher is better, $\downarrow$ lower is better). – indicates that not enough sample statistics were available. Units: Force in meV/Å, Stability in ps, $g_{A-A}$, $g_{A-B}$, $g_{B-B}$ in $10^{-3}$. Speed is listed in ns/d, based on a single 1000 ps simulation. Subscripts show standard deviations, if applicable. ”*” shown in Figure 1.
Lastly, we evaluate how the cutoff radius $r_{\text{cut}}$ and the correlation order $\nu$ affect the accuracy, computational performance, and stability of the MACE potential (Table 1). Overall, increasing either the correlation order or the cutoff radius reduces both RDF and force errors. [^1] Force errors can only be compared between models using the same mapping, as the mapping itself introduces an irreducible noise wang2019machine. A larger cutoff radius also substantially improves stability, though at the expense of higher computational cost. As expected, less expressive models run faster. As expected, reducing model complexity improves speed; however, classical models remain over an order of magnitude faster, achieving 1600, 2130, and 3470 ns/d for the four-, three-, and two-site models, respectively.
@@ 3.2 Capped Amino Acids
In recent decades, many different coarse-grained representations of amino acids have been developed for peptide and protein-related simulations. Mappings can range from near-atomistic representations that only remove hydrogens but preserve heavy-atom movements, to single bead mappings, in which each residue is represented by a single bead, for example the $C_{\alpha}$-atom kmiecik2016coarse. First, we discuss the widely studied capped L-alanine system (alanine dipeptide). This system has been extensively studied using different machine learning potentials thaler2022deep, thaler2023scalable, wang2019machine, wang2020ensemble, chen2025enhanced, as well as more recently generative models kohler2023flow, klein2025operator. We explore nine different mappings, ranging from an atomistic implicit solvent model to five-bead mappings, which preserve only the central backbone dihedrals. An overview of the mappings can be seen in Figure 3. For each mapping, we train a MACE model, and analyze the Ramachandran of the backbone dihdrals. We also perform 100 × 100 ps NVE simulations with different time steps to analyze stability (Supplementary Information).
@@
First, we discuss the results for high-resolution mappings in which solvent molecules and potentially hydrogen atoms are removed. Retaining only the heavy atoms, or a subset of them, is a common mapping choice, as it promises to preserve atomistic dynamics and makes comparisons with atomistic models straightforward.thaler2022deep, mirarchi2024amaro, wang2019machine, wang2020ensemble, charron2025navigating All high-resolution mappings are capable of accurately modeling the backbone dihedrals (Figure 3). In the NVE simulations, the implicit solvent model remains stable up to time steps of 1 fs, while the United and Heavy Atom model can be run safely up to 2 fs.
@@
The Core mapping, which miss the $C_{\beta}$ carbon, show a clear point symmetry around the center (Figure 3 column 2). This indicates that the CG molecule during the simulation freely switches between the L- and D-enantiomer. Since both the $C_{\alpha}$-hydrogen and the entire side chain are removed, the formal definition of chirality is lost, and a transition between enantiomers becomes an unhindered rotation of the backbone dihedrals. Adding the $C_{\beta}$ atom resolves this issue, and the CG molecule stays in the original L-enantiomer. We observed no noticeable improvement in terms of structural accuracy or stability when using the more detailed species assignment of Map 2. A more subtle symmetry was observed when using a single species for all beads. In this case, the model fails to capture the direction of the molecule, which results from the indistinguishability of local environments. We provide these findings, together with evaluations on the three other amino acids in the Supporting Information.
@@
Lastly, we show that while atomistic MLP models are chirally stable, CG models can undergo chiral inversion, i.e. the transition between enantiomers. To test this, we calculate the free energy barrier by biasing the improper $C_{\alpha}$-dihedral using well-tempered metadynamics. In all cases, we only train on L-alanine. We provide the exact biasing protocol in the Supporting Information.
@@
However, in CG representations, the transition becomes accessible. By removing one atom around the chiral center, in this case the $C_{\alpha}$-hydrogen, enatiomerization becomes accessible (Figure 4a). For the Heavy Atom and Core Beta mappings, we find a transition barrier of approximately 13 kcal/mol. The United Atom mapping presents a slightly larger barrier of 16 kcal/mol. This increase arises because merging heavy atoms with their hydrogens shifts the bead COM away from the stereocenter, thereby stabilizing the initial configuration. This also slightly shifts the minima of the improper dihedral.
@@ 3.3 Helix Formation of Polyalanine
As a last experiment, we extend the low resolution mappings from capped alanine to a capped alanine 15-mer. Polyalanine systems are good examples of peptides that spontaneously form alpha helices hazel2014thermodynamics, rudzinski2015bottom, kuczera2021length, zhou2007coarse, chakrabarti2001interrelationships. While naturally occurring L-amino acid peptides form right-handed helices (Figure 5a)sidorova2021protein, it has been experimentally shown that a D-alanine peptide forms the mirror image, a left-handed helix.milton1992
We tested three different mappings: the Core and Core Beta mapping from the capped alanine example, and a $C_{\alpha}$ mapping, which only preserves the $C_{\alpha}$ atom of each alanine residue. The $C_{\alpha}$ mapping is a popular choice for modeling large protein systems.tozzini2006mapping, yap2008coarse, kmiecik2016coarse, mioduszewski2023contact In the $C_{\alpha}$ mapping, we gave each $C_{\alpha}$ a different species. We also performed the experiment on three other possible choices: a single species, alternating species, and a symmetric species arrangement. Using unique or alternating species causes unphysical bond switches, similar to those in the two-site hexane model (Supporting Information).
Figure 5: (a) A chain of D-Alanine forms left-handed helices, while the naturally occurring L-Alanine forms right-handed helices. (b) Helicity and handedness of a 500 ns reference simulation of a capped 15-residue L-Alanine peptide.
Figure 6: Helicity and handedness of 100 × 5 ns trajectories based on different CG mappings. Left: $C_{\alpha}$, where each bead has a unique species embedding. Middle: Core. Right: Core Beta. The top row shows a graphical representation of the helix and per residue mapping.
@@ 4 Conclusions
We confirmed key findings by testing them with NequIP, a different E(3)-equivariant MLP.batzner20223 Our findings likely also extend to invariant graph neural networks, such as SchNetschutt2018schnet or DimeNetgasteiger2020directional, which rely on up to three-body scalar invariants (distances and angles). These models are blind to chirality, making them susceptible to the enantiomerization we observed.dumitrescu20243, adams2021learning Finally, bond permutations are likely to affect any architecture that constructs neighborhoods purely based on geometric information, without topological information.

### v2

@@ 1 Introduction
In this study, we investigate how modern E(3)-equivariant machine learning potentials can be applied to coarse-graining without the use of any priors. Using the MACE architecture as a representative modelMACE, we perform a systematic study to evaluate how the choice of mapping, species encoding and model parameters influence the stability and learned representation of the model. We compare these results with a classical CG potential to assess whether the learned models provide a more accurate description of coarse-grained energy landscapes. We evaluated three distinct systems of increasing complexity: liquid hexane, single amino acids, and a 15-mer polyalanine. We validate key findings by additionally testing them with NequIP batzner20223.
@@ 2.1 Coarse-graining
In coarse-graining (CG), the degrees of freedom of an atomistic system are reduced by grouping different atoms into beads. Here we limit the study to mappings, in which each atom at most contributes to one CG bead. We use the center-of-mass (COM) mapping to derive CG positions and forces. Other, more general mapping choicesbabadi2006coarse, gay1981modification, Voth2008-zb, as well as optimal mappingsdiggins2018optimal, yang2023slicing, wang2019coarse, have also been explored, but remain less prevalent in practice. In the COM mapping, the CG positions $\textbf{R}\in\mathcal{R}^{3N}$ are derived as a linear mapping M of the atomistic positions $\textbf{r}\in\mathcal{R}^{3n}$, where $N<n$:
@@ 2.3 Coarse-grained Potential
To parameterize an equivariant MLP $U_{MLP}(\mathbf{R};\theta)$, we employ the MACE architectureMACE. MACE combines the idea of atomic cluster expansiondrautz2019atomic with message-passing neural networks (MPNNs)gilmer2017neural. In MACE, the messages passed between nodes, in our case CG beads, are expanded using the idea of a hierarchical body order expansion over neighbors $J$:
@@ 2.7 Well-tempered Metadynamics
Here $h$ is the initial Gaussian height, $\sigma_{i}$ is the width of the $i$-th collective variable, and $\gamma=T^{*}/T$ is the bias factor. This corresponds to sampling the collective variables at an effective temperature $T^{*}=\gamma T$, which scales the free energy barriers by $1/\gamma$.
@@ 3.1 Liquid Hexane
Table 1: Comparison of a 2-layer MACE model with different correlation orders $\nu$ and average number of neighbors $N_{avg}$. Results are based on 50 × 1000 ps simulations. Green highlighting indicates better performance (arrows indicate direction: $\uparrow$ higher is better, $\downarrow$ lower is better). – indicates that not enough sample statistics were available. Units: Force in meV/Å, stability in ps, $g_{A-A}$, $g_{A-B}$, $g_{B-B}$ in $10^{-3}$. Speed is listed in ns/d, based on a single 1000 ps simulation. Subscripts show standard deviations, if applicable. * signifies the model shown in Figure 1.
Lastly, we evaluate how the cutoff radius $r_{\text{cut}}$ and the correlation order $\nu$ affect the accuracy, computational performance, and stability of the MACE potential (Table 1). Overall, increasing either the correlation order or the cutoff radius reduces both RDF and force errors. [^1] Force errors can only be compared between models using the same mapping, as the mapping itself introduces an irreducible noise wang2019machine. A larger cutoff radius also substantially improves stability, though at the expense of higher computational cost. As expected, less expressive models run faster. Similarly, reducing model complexity improves speed; however, classical models remain over an order of magnitude faster, achieving around 1600, 2130, and 3470 ns/d for the four-, three-, and two-site models, respectively.
@@ 3.2 Capped Amino Acids
In recent decades, many different coarse-grained representations of amino acids have been developed for peptide and protein-related simulations. Mappings can range from near-atomistic representations that only remove hydrogens but preserve heavy-atom movements, to single bead mappings, in which each residue is represented by one bead, for example the $C_{\alpha}$-atom kmiecik2016coarse. First, we discuss the widely studied capped L-alanine system (alanine dipeptide). This system has been extensively studied using different machine learning potentials thaler2022deep, thaler2023scalable, wang2019machine, wang2020ensemble, chen2025enhanced, as well as more recently generative models kohler2023flow, klein2025operator. We explore nine different mappings, ranging from an atomistic implicit solvent model to five-bead mappings, which preserve only the central backbone dihedrals. An overview of the mappings can be seen in Figure 3. For each mapping, we train a MACE model, and analyze the Ramachandran of the backbone dihdrals. We also perform 100 × 100 ps NVE simulations with different time steps to analyze simulation stability (Supplementary Information).
@@
First, we discuss the results for high-resolution mappings in which solvent molecules and potentially hydrogen atoms are removed. Retaining only the heavy atoms, or a subset of them, is a common mapping choice, as it promises to preserve atomistic dynamics and makes comparisons with atomistic models straightforward.thaler2022deep, mirarchi2024amaro, wang2019machine, wang2020ensemble, charron2025navigating All high-resolution mappings are capable of accurately modeling the backbone dihedrals (Figure 3). In the NVE simulations, the implicit solvent model remains stable up to time steps of 0.5 fs, while the United and Heavy Atom model can be run safely up to 3 fs.
@@
The Core mappings, which miss the $C_{\beta}$ carbon, show a clear point symmetry around the center (Figure 3 column 2). This indicates that the CG molecule during the simulation freely switches between the L- and D-enantiomer. Since both the $C_{\alpha}$-hydrogen and the entire side chain are removed, the formal definition of chirality is lost, and a transition between enantiomers becomes an unhindered rotation of the backbone dihedrals. Adding the $C_{\beta}$ atom resolves this issue, and the CG molecule stays in the original L-enantiomer. We observed no noticeable improvement in terms of structural accuracy or stability when using the more detailed species assignment of Map 2. A more subtle symmetry was observed when using a single species for all beads. In this case, the model fails to capture the direction of the molecule, which results from an indistinguishability of local environments. We provide these findings, together with evaluations on three other amino acids in the Supporting Information.
@@
Lastly, we show that CG amino acid models can undergo chiral inversion, i.e. they transition between enantiomers, even if only a single atom is removed around the stereocenter. For that, we calculate the free energy barrier for the transition by biasing the improper $C_{\alpha}$-dihedral using well-tempered metadynamics. In all cases, we only train on L-alanine. We provide the exact biasing protocol in the Supporting Information.
@@
By removing one atom around the chiral center, in this case the $C_{\alpha}$-hydrogen, enantiomerization becomes accessible (Figure 4a). For the Heavy Atom and Core Beta mappings, we find a transition barrier of approximately 13 kcal/mol. The United Atom mapping presents a slightly larger barrier of 16 kcal/mol. This increase arises because merging heavy atoms with their hydrogens shifts the bead COM away from the stereocenter, thereby stabilizing the initial configuration. This also slightly shifts the minima of the improper dihedral.
@@ 3.3 Helix Formation of Polyalanine
As a last experiment, we extend the low resolution mappings from capped alanine to a capped 15-mer alanine. Polyalanine systems are good examples of peptides that spontaneously form alpha helices hazel2014thermodynamics, rudzinski2015bottom, kuczera2021length, zhou2007coarse, chakrabarti2001interrelationships. While naturally occurring L-amino acid peptides form right-handed helices (Figure 5a)sidorova2021protein, it has been experimentally shown that a D-alanine peptide forms the mirror image, a left-handed helix.milton1992
We tested three different mappings: the Core and Core Beta mapping from the capped alanine example, and a $C_{\alpha}$ mapping, which only preserves the $C_{\alpha}$ atom of each alanine residue. The $C_{\alpha}$ mapping is a popular choice for modeling large protein systems.tozzini2006mapping, yap2008coarse, kmiecik2016coarse, mioduszewski2023contact In the $C_{\alpha}$ mapping, we gave each $C_{\alpha}$ a different species. We also performed experiments on three other possible choices: a single species, alternating species, and a symmetric species arrangement. Using unique or alternating species causes unphysical bond switches and ultimately instabilities. We all the results of all $C_{\alpha}$ mappings in the Supporting Information.
Figure 5: (a) A chain of D-Alanine forms left-handed helices, while the naturally occurring L-Alanine forms right-handed helices. (b) Helicity and handedness of a 500 ns reference simulation of a capped 15-mer L-alanine.
Figure 6: Helicity and handedness of 100 × 5 ns trajectories based on different CG mappings. Left: $C_{\alpha}$, where each bead has a unique species embedding. Middle: Core. Right: Core Beta. The top row shows a graphical representation of the ideal helix and per residue mapping.
@@ 4 Conclusions
We confirmed key findings with NequIP, a different E(3)-equivariant MLP.batzner20223 Our findings likely also extend to invariant graph neural networks, such as SchNetschutt2018schnet or DimeNetgasteiger2020directional, which rely on up to three-body scalar invariants (distances and angles). These models are blind to chirality, making them susceptible to the enantiomerization we observed.dumitrescu20243, adams2021learning Finally, bond permutations are likely to affect any architecture that constructs neighborhoods purely based on geometric information, without topological information.
@@
5 Data and Software Availability
The code and data supporting this study are publicly available at https://github.com/tummfm/CG-Mapping-Benchmark. The training framework chemtrain is publicly available at https://github.com/tummfm/chemtrain.